Direction: In each of the following number series questions one number is wrong. You have to identify that number
35 58 87 119 155 196

Answer by Edwin McCravy(20056)

35   58   87  119  155  196

Let's form a difference table.  To do that let's 
subtract each from its right neighbor and place
the diference between and below wach pair subtracted
on the second row. Then do the same with that row 
and the row below it.

35   58   87  119  155  196
   23   29   32   36   41
      6    3    4    5 
        -3    1    1

There are 2 1's on the bottom row.  Let's change the
-3 on the bottom row to 1 also, so we'd have all 1's 
on the bottom row.

35   58   87  119  155  196
   23   29   32   36   41
      6    3    4    5 
         1    1    1

That would make the 6 have to be a 2, so that 3-2 = 1
So let's change the 6 to 2

35   58   87  119  155  196
   23   29   32   36   41
      2    3    4    5 
         1    1    1

That would make the 23 have to be a 27, so that 29-27=2
So let's change the 23 to 27

35   58   87  119  155  196
   27   29   32   36   41
      2    3    4    5 
         1    1    1

That would make the 35 have to be a 31, so that 58-31=27.
So let's change the 35 to 31

31   58   87  119  155  196
   27   29   32   36   41
      2    3    4    5 
         1    1    1

Answer: the first number 35, is wrong, and should be 31.

Edwin